OpenStudy (anonymous):

At 2:00 p.m. two cars start toward each other from towns 240 miles apart if the rate of one car is 10mph faster than the other, how fast does each car go if they meet at 5:00 P.M. If 3x represents the distance of the slower car traveled then which expression represents the distance the faster car traveled

OpenStudy (anonymous):

(x+10)*3+3*x=240 x=35 x+10=45

OpenStudy (anonymous):

slower car rate = 105/3 = 35 mph faster car rate = 35+10 = 45 mph